# Exploring Scene Coherence for Semi-Supervised 3D Semantic Segmentation

### v1

@@ Introduction
Table 2: Ablation study of different components of our designed CoScene. PtErase refers to the point erause strategy, while MixPatch and InsFill refers to the patch mix and instance filling data augmentation, respectively. All mIoU results are reported in percentage (%).
@@
Figure 5: Performance comparison on SemanticKITTI val split under different semi-supervised settings. (a) Illustration of performance differences whether using consistency assurance strategy on the LaserMix (Kong et al. 2023b). (b) Comparison between dual-scene level and multi-scene level for the patch-based data augmentation in our CoScene network. (c) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
Effectiveness of point erasure strategy. The row 1 in Table 2 denotes the performance of MeanTeacher (Tarvainen and Valpola 2017) as the baseline, revealing a marginal enhancement compared to the supervision-only Cylinder3D model (Zhu et al. 2021). When PtErase is applied to the baseline, it is effective and can boost the performance by ensuring consistency during forward and backward propagation, thereby compensating for the shortcomings of the pseudo-label mechanism. The consistency assurance is implemented through a plug-and-play point erasure strategy, which can be easily integrated into methods based on the pseudo-label mechanism. To further validate the significance of ensuring consistency during forward and backward propagation, we embedded the PtErase strategy into LaserMix. As illustrated in Figure 5 (a), performance improves across all ratios when using the PtErase strategy. This further emphasizes the importance of intra-scene consistency and demonstrates the generalization ability of our method.
@@
Effectiveness of patch-based data augmentation. The patch-based data augmentation consists of MixPatch (patch-based mix) and InsFill (instance filling). Rows 3 and 4 of Table 2 demonstrate that using MixPatch or InsFill alone results in an average mIoU improvement of 3.1% or 3.4%, respectively, compared to the baseline. This suggests that annotation information transfer across scenes benefits SSL 3D segmentation. As shown in rows 5 and 6 of the Table 2, combining MixPatch or InsFill with PtErase significantly enhances performance, which indicates the effectiveness of the patch-based data augmentation. Furthermore, the collaboration of all three components, leads to substantial performance gains, as validated by the last row of Table 2. Additionally, our patch pool and instance pool within the patch-based data augmentation facilitate effective information fusion across multiple scenes. This approach inherently enables the transfer of inter-scene diverse correlation information. To verify the effectiveness of the interaction at the multi-scene level, we modify the pool implementation to enable only interaction between two scenes. The comparative results, as shown in Figure 5 (b), indicate superior performance at multi-scene level compared to dual-scene level.
@@
Different mixing strategies. Figure 5 (c) compares the patch-based data augmentation employed in our CoScene with other mixing approaches (Nekrasov et al. 2021; Xiao et al. 2022; Yun et al. 2019; Zhang et al. 2017). MixUp demonstrates a slight improvement over the baseline, likely due to its limited ability to transform information. Other methods have a considerable improvement over the baseline by enriching point cloud distributions. In contrast, our CoScene achieves the greatest gain by effectively exploring rich inter-scene correlation, highlighting its effectiveness in facilitating information transfer between labeled and unlabeled scenes. This is accomplished through its innovative design, which includes a patch pool and an instance pool spanning multiple scenes.
@@
Patch split number. As the patch split number $n$ plays a critical role in our framework, we further analyze the impact of varying $n$ in patch-based data augmentation, with the results presented in Figure 6. When $n$ is small, it can yield a certain level of enhancement, but it may fail to generate data sufficiently divergent from the original, impeding the effective transfer of information between labeled and unlabeled scenes. Conversely, when $n$ is too large, the generated scenes may deviate excessively from the original data, resulting in the loss of essential prior scene information. Empirically, we identify that setting $n$ to 18 yields optimal results. Therefore, throughout this paper, we consistently use $n$ set to 18, unless explicitly stated otherwise.
Figure 6: Comparison of different mixing-based techniques.
Conclusion
In this paper, we explore the scene coherence for semi-supervised point cloud segmentation, building upon the intra-scene consistency and inter-scene correlation. In particular, we achieve intra-scene consistency assurance through a point erasure strategy that ensures consistency during the forward and backward propagation, rather than simply disregarding points with pseudo-labels as in previous methods. Additionally, we explore inter-scene correlation and propose a patch-based data augmentation to enhance information transfer between labeled scenes and unlabeled scenes. Benefiting from these two components, our CoScene effectively transfers diverse semantic information and produces more accurate segmentation results. Extensive experimental results demonstrate the superiority of our CoScene over state-of-the-art methods.

### v2

@@ 4.4 Ablation Study
Table 3: Ablation study of different components of our AIScene framework. PtErase refers to the point erasure strategy, while MixPatch and InsFill refer to the patch mix and instance filling data augmentation, respectively.
[55]
[49]
[48]
[26]
[45]
[15]
[5]
@@
Figure 6: (a) Illustration of performance differences whether using point erasure strategy on the LaserMix [15]. (b) Variation in the proportion of points removed by the point erasure strategy throughout the training process. (c) Comparison of different mixing-based techniques. (d) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
Figure 7: The left chart shows the initial output with distributed probabilities across multiple classes, indicating uncertainty in class prediction. Through our point erasure strategy, the right chart exhibits a distribution with a dominant class (“person”) having a significantly higher probability,
@@
Effectiveness of Patch-based Data Augmentation. The patch-based data augmentation consists of MixPatch (patch-based mixing) and InsFill (instance filling). In Tab. 3, the results in the 3rd and 4th rows indicate that using MixPatch or InsFill individually yields an average mIoU improvement of 4.0% or 4.3%, respectively, over the baseline. This suggests that incorporating diverse semantic information across scenes benefits semi-supervised LiDAR segmentation. Furthermore, as shown in the 5th and 6th rows, combining MixPatch or InsFill with PtErase significantly enhances performance, highlighting the effectiveness of patch-based data augmentation. The combination of all three components yields substantial performance gains, as evidenced by the last row. Additionally, the patch pool and instance pool within the patch-based data augmentation facilitate the fusion of diverse semantic information across multiple scenes. To verify the effectiveness of interactions at the multi-scene level, we modify the pool implementation to allow interactions between only two scenes. The results in Tab. 4 indicate that multi-scene interactions yield superior performance compared to dual-scene interactions.
@@
Different Mixing Strategies. Fig. 6 (c) compares our patch-based data augmentation with existing mixing methods [26, 45, 48]. MixUp [49] shows a slight improvement over the baseline, likely due to its limited capacity for information transformation. Note that we reproduce CylinderMix in the LiDAR branch [5] for comparison. Other methods demonstrate considerable improvement over the baseline by enriching point cloud distributions. In contrast, our AIScene achieves the most substantial gains by effectively leveraging inter-scene correlation, highlighting its effectiveness in facilitating diverse semantic information generation for labeled and unlabeled scenes. This is achieved through an innovative design that incorporates a patch pool and an instance pool spanning multiple scenes.
Splits on Semantic KITTI
Augmentation Type Methods 1% 10% 20% 50% Avg.
dual-scene level Ours 53.4 61.3 61.8 62.2 59.7
multi-scene level Ours 54.5 63.3 63.7 64.3 61.5
Improvement - +1.1 +2.0 +1.9 +2.1 +1.8
Table 4: Comparison between dual-scene and multi-scene levels in the patch-based data augmentation of our AIScene.
@@
Patch Split Number. As the patch split number $n$ plays a critical role in our framework, we further analyze the impact of varying $n$ in patch-based data augmentation, with the results presented in Fig. 6 (d). When $n$ is small, it can yield a certain level of enhancement, but it may fail to generate data sufficiently divergent from the original, thereby impeding the effective generation of diverse semantic information between labeled and unlabeled scenes. Conversely, when $n$ is too large, the generated scenes may deviate excessively from the original data, resulting in a loss of essential prior scene information. Empirically, we find that setting $n$ to 18 yields optimal results. Throughout this paper, we consistently use $n$ set to 18 unless explicitly stated otherwise.
5 Conclusion
This paper explores scene affinity for semi-supervised LiDAR semantic segmentation in driving scenes, named AIScene. Specifically, AIScene addresses intra-scene inconsistency through a point erasure strategy, which enhances the learning of pseudo-labeled points. Moreover, it employs a patch-based data augmentation method to mix multiple scenes at both scene and instance levels, guided by inter-scene correlation. This method enriches the semantic diversity of labeled scenes and complements pseudo-labeled scenes. Extensive experiments show that AIScene achieves state-of-the-art results on popular benchmarks, with remarkable improvements in the most challenging settings. Additionally, the novel erasure strategy and augmentation method can function independently to facilitate semi-supervised LiDAR semantic segmentation.